how much water should be added to 12 mL of 17% alcohol solution to reduce the concentration to 6%?
Answer by mananth(16949)

12 mL of 17% alcohol solution
mix concentration 6% 12+x ml
water 0% x ml
..
0+12*0.17=0.06(12+x)
2.04 = 0.72 +0.06x
add -0.72 to both sides
2.04 - 0.72 = 0.72-0.72+0.06x
1.38=0.06x
divide by 0.06
1.38/0.06 = x
23 =x
23 ml water